40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 18:42:43 2014, Jenne, Update, LSC, Green PDH box boosts BoostCartoon.pdf
    Reply  Thu Sep 4 00:56:44 2014, Jenne, Update, LSC, Green PDH box boosts TFstage_newDesign_3Sept2014.pngBoostNoBoost_newDesign_3Sept2014.pngBoostOnly_newDesign_3Sept2014.pngD0901351-v1_3Sept2014.pdf
       Reply  Thu Sep 4 03:12:55 2014, ericq, Update, LSC, Green PDH box boosts boostedTF.pdfgain.jpg
       Reply  Thu Sep 4 16:45:10 2014, Jenne, Update, LSC, Green PDH box boosts IMG_1712.JPGIMG_1714.JPG
          Reply  Fri Sep 5 05:32:57 2014, Jenne, Update, LSC, Green PDH box boosts TRX_TRY_ALSvsIR_4Sept2014.pdf
             Reply  Fri Sep 5 21:13:57 2014, Jenne, Update, LSC, Green PDH out of loop ALS_XY_outOfLoop_5Sept2014.pdf
Message ID: 10452     Entry time: Thu Sep 4 16:45:10 2014     In reply to: 10448     Reply to this: 10458
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Green PDH box boosts 

As EricQ mentioned in last night's elog, the modifications were made to the Yend (SN 17) uPDH board.

R31 became 49.9 Ohms, R30 became 45.3kOhm, R24 became 1.02k, R16 became 1k, a new flying resistor is tombstoned up against R24 and connected by purple wire to C6 and it is 20k.  C28 is 183nF and C6 is 100nF.  These numbers were used in Q's simulation last night.

 

 

IMG_1712.JPGIMG_1714.JPG

ELOG V3.1.3-